INAF queue observing, 2015 Sep 30-Oct 01

Observer: F. Cusano, R. Carini
Telescope Operator: Geno Bechetti
Support Astronomer: Dave Thompson
Instrument: LBC, MODS


At the sunset the sky is mostly clear, there are some cirrus, humidity 50%

Night duration (12\xB0 el twilight) 19:30-4:50 ( 02:30-11:50 UT) = 9 h 20m

Observing time: 8 hrs 50 m

Weather time loss: 0 hrs (0%)

technical time loss 30 m

Data Summary

At the sunset the sky is clear, humidity 50%.

Project_RUN Objects Filter ExpTime Notes
Nascimbeni Wasp 52 dual grating+ lbc red 4 h 30m COMPLETED
standard G191-B2B dual and red grating    
Antoniucci v512per ima+ dual grating 25m COMPLETED
Antoniucci v1118ori ima+ dual grating 25m COMPLETED
Antoniucci SDSSJ0510 ima + dual grating 25m COMPLETED


02:26 start Nascimbeni program, we acquire the target Wasp 52
  • acqMODS mods.4.wasp52.acq mods1r.20151001.0002.fits-0004
  • 02:40 primary mirror panic

02:50 we start the acquisition again mods1r.0005-0007
  • there are problems with the primary mirror.

03:02 collimation LBC: dofpia, /first, /X2

03:15 begin the science
  • run the LBC ob: wasp52_f972.ob, LBC should be defocused until the most brightest star among WASP-52 and REF#1 (see chart) has a peak value of about ~30,000 ADU, as required by the PI
  • execMODS mods.4.wasp52.obs mods1r.0008-0087, mods1b.0002-0048
  • 03:30 found the right defocus, run wasp52_f972.ob
  • seeing 1.0''
  • 07:17 seeing 0.8''

07:48 slew to standard star G191-B2B
  • acqMODS g191b2b.acq mods1r.0088-0091
  • execMODS g191b2b.obs mods1r.0092-0094,mods1b.0049-0051 dual grating
  • mods1r.0095-0097 red grating

08:27 slew to Antoniucci's program, target v512per
  • 08:30 the TO unwrap
  • 08:37 execMODS v512per.img mods1r.0098-0100,mods1b.0052-0054
  • 08:48acqMODS v512per_ls.acq mods1r.0101-0105
  • 09:00 execMODS v512per.obs mods1r.0106-0109,mods1b.0055-0058

09:40 slew to the target V1118Ori
  • execMODS v1118ori.img mods1r.0110-0113,mods1b.0059-0062
  • mods1b.0060 and mods1r.0111 are not good
  • seeing 0.9''
  • 09:55 acqMODS v1118ori_ls.acq mods1r.0114-0116
  • 10:03 execMODS v1118ori_ls.obs mods1r.0117-0120,mods1b.0063-0066
  • 10:15 seeing 0.6''

10:34 slew to the targets SDSSJ0510
  • execMODS sdssj0510.img mods1r.0121-0123-,mods1b.0067-0069 are saturated
  • seeing 0.7''
  • decrease the time exposure to 10 sec mods1r.0124-0126,mods1b.0070-0072
  • 10:50 acqMODS sdssj0510.acq mods1r.0127-0129
  • 11:00 execMODS sdssj0510.obs mods1r.0130-0133,mods1b.0073-0076

11:26 slew to target v2775ori
  • execMODS v2775.img mods1r.0134-0136,mods1b.0077-0079 the target is not visible
  • 11:38 acqMODS v2775ori.acq mods1r.0137 the target is not visible
  • increase exp time to 80 s mods1r.0138 the target is not visible
  • increase to 130 sed mods1r.0139 the target is not visible
  • increase to 200 sec mods1r.0140 the target is not visible

00:00 flat field LBC, are not good, problems with the dithering

01:00 bias LBC 25BIAS_BINO
  • calibration for the target of Nascimbeni execMODS (arc + flat field) mods1r.0141-0142, mods1b.0080-0081 are not good, problems with the dashboard, it was stucked.
  • execMODS again mods1r.0143-0147, mods1b.0082-0088
  • dual grating slitless flat execMODS mods1r.0148-0152,mods1b.0089-0096
  • BIAS MODS dobias mods1r.0153-0162, mods1b.0097-0105
  • calibration for the standard star. flat filed with slit 5'' mods1r.0163-0165, mods1b.0106-0111
  • arc for the standard star mods1r.0166, mods1b.0112-0115

-- %USERSIG{RobertaCarini - 2015-10-01}%


Topic revision: r11 - 26 Nov 2015, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback